YoursTruly:

Add and simplify: (9x + 3) + (2x + 5) A. 19+x B. 8x+11 C. 19x D. 11x+8

iGreen:

9x + 3 + 2x + 5 Add like terms 9x + 2x = ? 3 + 5 = ?

AMQ:

lmao ^

AMQ:

just the way you did it on openstudy, huh?

iYuko:

He did it already......

Warriorz13:

T

Warriorz13:

This is a really easy equation.

YoursTruly:

What the ***... I just posted that question. You can't have an answer already. .-.

iGreen:

Adding terms with 'x' is the same as basic addition. Just add the constants and keep the 'x'.

iYuko:

iGreen Best Response Undo Medals 1 9x + 3 + 2x + 5 Add like terms 9x + 2x = ? 3 + 5 = ?

AMQ:

answer is D

AMQ:

xd

YoursTruly:

Dude, I literally posted this question not even a minute ago... how tf you gonna get an answer that quick.

YoursTruly:

I'm giving AMQ Best Response because something ain't right about iGreen...

AMQ:

;)

iGreen:

I'm a robot.

YoursTruly:

I got a 100% o;;;

AMQ:

yw ;)

YoursTruly:

Okay, now I believe that. I believe iGreen is a robot. x.x

AMQ:

nah

AMQ:

he's just a math ***head

AMQ:

*dic'k head

YoursTruly:

x'D

Bob:

@iGreen is an old user from OS.

Bob:

Get your censoring bypasser on. @AMQ

AMQ:

yeah, it's currently not working, soooo.... :(